. Water from an irrigation sprinkler lands on a corn plant 37 meters away. If the sprinkler angle is 40˚,
with what initial velocity was the water launched?
1
Expert's answer
2020-09-06T17:17:16-0400

Initially, water velocity had two components, y and x. The range (37 m) is defined by both:


"v_y=v\text{ sin}\theta,\\
v_x=v\text{ cos}\theta."

The time required for a mass of water to go up:


"t_\text{up}=\frac{v_y}{g}."

It is easy to understand that the time for the water to reach the ground level again is

"t=2t_\text{up}=2\frac{v_y}{g}=\frac{2v\text{ sin}\theta}{g}."

The range is determined by the horizontal component only:


"R=v_xt=v\text{ cos}\theta\cdot\frac{2v\text{ sin}\theta}{g}=\frac{v^2\text{sin}(2\theta)}{g}."

Express and calculate the initial velocity:


"v=\sqrt{\frac{gR}{\text{sin}(2\theta)}}=19.2\text{ m/s}."